# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ Appendix C Link with Cauchy’s steepest descent and Newton’s method
$\bm{\theta}_{t+1}:=\bm{\theta}_{t}-\mathbf{G}_{t}\mathds{1}_{S}(\mathds{1}_{S}^{T}\mathbf{G}_{t}\mathbf{H}_{t}\mathbf{G}_{t}\mathds{1}_{S})^{-1}\mathds{1}_{S}^{T}\mathbf{G}_{t}\mathbf{g}_{t}=\bm{\theta}_{t}-\mathbf{g}_{t}\frac{\mathbf{g}_{t}^{T}\mathbf{g}_{t}}{\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}},$ (32)
@@
$\bm{\theta}_{t+1}$ $:=\bm{\theta}_{t}-\mathbf{G}_{t}(\mathbf{G}_{t}\mathbf{H}_{t}\mathbf{G}_{t})^{-1}\mathbf{G}_{t}\mathbf{g}_{t}.$ (33)
@@
$\bm{\theta}_{t+1}$ $:=\bm{\theta}_{t}-\mathbf{G}_{t}\mathbf{x}=\bm{\theta}_{t}-\mathbf{H}_{t}^{-1}\mathbf{g}_{t},$ (35)
@@
Appendix D Anisotropic Nesterov cubic regularization
@@
$T(\bm{\eta}):=-\bm{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\bm{\eta}\bar{\mathbf{H}}\bm{\eta}+\frac{\lambda_{\mathrm{int}}}{6}\|\mathbf{D}\bm{\eta}\|^{3}.$ (36)
The function $T$ is strictly convex if, and only if, $\bar{\mathbf{H}}$ is positive definite. Moreover, $T$ is differentiable twice and has at least one global minimum $\bm{\eta}_{*}$, so $\frac{\mathrm{d}T}{\mathrm{d}\bm{\eta}}(\bm{\eta}_{*})=0$. Therefore, we first look for the solutions of the equation $\frac{\mathrm{d}T}{\mathrm{d}\bm{\eta}}(\bm{\eta})=0$.
@@
$\frac{\mathrm{d}T}{\mathrm{d}\bm{\eta}}(\bm{\eta})$ $=-\bar{\mathbf{g}}+\bar{\mathbf{H}}\bm{\eta}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\bm{\eta}\|\mathbf{D}^{2}\bm{\eta}$
$=-\bar{\mathbf{g}}+\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\bm{\eta}\|\mathbf{D}^{2}\right)\bm{\eta},$
@@
$\bar{\mathbf{g}}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\bm{\eta}\|\mathbf{D}^{2}\right)\bm{\eta}.$ (37)
Let $\bm{\eta}^{\prime}:=\mathbf{D}\bm{\eta}$. Eqn. 37 is then equivalent to:
$\bar{\mathbf{g}}$ $=\left(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}\|\bm{\eta}^{\prime}\|\mathbf{D}\right)\bm{\eta}^{\prime}.$
$=\frac{\lambda_{\mathrm{int}}}{2}\mathbf{D}\left(\frac{2}{\lambda_{\mathrm{int}}}\mathbf{D}^{-1}\bar{\mathbf{H}}\mathbf{D}^{-1}+\|\bm{\eta}^{\prime}\|\mathbf{I}\right)\bm{\eta}^{\prime}$
@@
$\bar{\mathbf{g}}$ $=\frac{\lambda_{\mathrm{int}}}{2}\mathbf{D}\left(\mathbf{K}+\|\bm{\eta}^{\prime}\|\mathbf{I}\right)\bm{\eta}^{\prime}$ (38)
@@
In this case, Eqn. 38 is equivalent to:
$\bm{\eta}^{\prime}$ $=\frac{2}{\lambda_{\mathrm{int}}}\left(\mathbf{K}+\|\bm{\eta}^{\prime}\|\mathbf{I}\right)^{-1}\mathbf{D}^{-1}\bar{\mathbf{g}}.$
Now, let $r=\|\bm{\eta}^{\prime}\|$. We want to solve:
@@
Trivially: $\bm{\eta}$ solution of 37 $\Rightarrow$ $\mathbf{D}\bm{\eta}$ solution of 38 $\Rightarrow$ $\|\mathbf{D}\bm{\eta}\|$ solution of 39. Reciprocally: $r$ solution of 39 $\Rightarrow$ $\bm{\eta}^{\prime}:=(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}r\mathbf{D})^{-1}\bar{\mathbf{g}}$ solution of 38 $\Rightarrow$ $\mathbf{D}^{-1}\bm{\eta}^{\prime}$ solution of 37.
Therefore, in order to find the unique global minimum of $T$, it is sufficient to solve Eqn. 39. This is doable numerically.
@@
We follow the procedure proposed in (Nesterov & Polyak, 2006, Section 5). Let $\lambda_{\mathrm{min}}$ be the minimum eigenvalue of $\mathbf{K}$. So, $\lambda_{\mathrm{min}}\leq 0$. Following Nesterov & Polyak (2006), we look for the unique $\bm{\eta}^{\prime}$ belonging to $\mathcal{C}:=\{\bm{\eta}^{\prime}\in\mathbb{R}^{S}:\|\bm{\eta}^{\prime}\|>|\lambda_{\mathrm{min}}|\}$, which is also the solution of maximum norm of Eqn. 38. Conditionally to $\bm{\eta}^{\prime}\in\mathcal{C}$, $(\mathbf{K}+\|\bm{\eta}^{\prime}\|\mathbf{I})$ is invertible. So we only need to solve:
@@
which has exactly one solution $r_{*}$. Then, we compute $\bm{\eta}_{*}:=\mathbf{D}^{-1}(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}r_{*}\mathbf{D})^{-1}\bar{\mathbf{g}}$.
Appendix E Invariance by subset-wise affine reparameterization
E.1 Motivation
The choice of the best per-layer parameterization is still a debated question. On the theoretical side, the standard parameterization cannot be used to train very wide networks, because it leads to a diverging first gradient step Yang & Hu (2021). Besides, the NTK parameterization is widely used in theoretical works in order to manage the infinite-width limit Jacot et al. (2018); Du et al. (2019); Arora et al. (2019); Lee et al. (2019); Mei & Montanari (2022). On the practical side, the standard parameterization is preferred over the NTK one because it leads to better results, both in terms of training and generalization.
@@ E.1 Motivation
E.2 Claim
We consider a parameter $\tilde{\bm{\theta}}$ such that $\bm{\theta}=\varphi(\tilde{\bm{\theta}})$, where $\varphi$ is an invertible map, affine on each subset of parameters. Therefore, its Jacobian is: $\mathbf{J}=\mathrm{Diag}(\alpha_{1},\cdots,\alpha_{p})$, where, for all $1\leq s\leq S$ and $1\leq p_{1},p_{2}\leq P$, we have:
@@ E.2 Claim
We want to compare the training trajectory of $\mathcal{L}(\bm{\theta})$ and $\mathcal{L}(\varphi(\tilde{\bm{\theta}}))$ when using Method 1. For any quantity $\mathbf{x}$ computed with the parameterization $\bm{\theta}$, we denote by $\tilde{\mathbf{x}}$ its counterpart computed with the parameterization $\tilde{\bm{\theta}}$.
We compute $\tilde{\bm{\eta}}_{*}$. Equation 10 gives:
$\tilde{\bm{\eta}}_{*}=\left(\tilde{\bar{\mathbf{H}}}+\frac{\lambda_{\mathrm{int}}}{2}\|\tilde{\mathbf{D}}\tilde{\bm{\eta}}_{*}\|\tilde{\mathbf{D}}^{2}\right)^{-1}\tilde{\bar{\mathbf{g}}}.$ (42)
@@
since $\mathbf{J}$ and $\mathbf{U}$ are diagonal. And, since $\mathbf{D}_{ii}=|(\mathbf{D}_{\bm{\theta}}^{(3)}(\mathbf{u}))_{iii}|^{1/3}$, then $\tilde{\mathbf{D}}_{ii}=a_{i}^{2}\mathbf{D}_{ii}$, thus $\tilde{\mathbf{D}}=\tilde{\mathbf{J}}^{2}\mathbf{D}$.
Thus, Eqn. 42 becomes:
$\tilde{\bm{\eta}}_{*}=\left(\tilde{\mathbf{J}}^{2}\bar{\mathbf{H}}\tilde{\mathbf{J}}^{2}+\frac{\lambda_{\mathrm{int}}}{2}\|\tilde{\mathbf{J}}^{2}\mathbf{D}\tilde{\bm{\eta}}_{*}\|\tilde{\mathbf{J}}^{4}\mathbf{D}^{2}\right)^{-1}\tilde{\mathbf{J}}^{2}\bar{\mathbf{g}},$
@@
$\tilde{\mathbf{J}}^{2}\tilde{\bm{\eta}}_{*}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\tilde{\mathbf{J}}^{2}\tilde{\bm{\eta}}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}}.$
Therefore, $\tilde{\bm{\eta}}_{*}$ is a solution of Eqn. 10 in the parameterization $\tilde{\bm{\theta}}$ if, and only if, $\tilde{\mathbf{J}}^{2}\tilde{\bm{\eta}}_{*}$ is a solution in the parameterization $\bm{\theta}$. Moreover, $\|\tilde{\mathbf{D}}\tilde{\bm{\eta}}_{*}\|=\|\mathbf{D}\mathbf{J}^{2}\tilde{\bm{\eta}}_{*}\|$, so $\tilde{\bm{\eta}}_{*}$ is the solution of maximum norm $\|\tilde{\mathbf{D}}\tilde{\bm{\eta}}_{*}\|$ of 10 with parameterization $\tilde{\bm{\theta}}$ iff $\tilde{\mathbf{J}}^{2}\tilde{\bm{\eta}}_{*}$ is a the solution of maximum norm $\|\mathbf{D}\mathbf{J}^{2}\tilde{\bm{\eta}}_{*}\|$ of 10 with parameterization $\bm{\theta}$.
Thus, $\bm{\eta}_{*}=\tilde{\mathbf{J}}^{2}\tilde{\bm{\eta}}_{*}$, and the update step in parameterization $\tilde{\bm{\theta}}$ is:
$\tilde{\bm{\theta}}_{t+1}$ $=\tilde{\bm{\theta}}_{t}-\tilde{\mathbf{U}}_{t}\mathbf{I}_{P:S}\tilde{\bm{\eta}}_{*}$
$=\tilde{\bm{\theta}}_{t}-\tilde{\mathbf{U}}_{t}\mathbf{I}_{P:S}\tilde{\mathbf{J}}^{-2}\bm{\eta}_{*},$
@@
$\mathbf{J}^{-1}\bm{\theta}_{t+1}$ $=\mathbf{J}^{-1}\bm{\theta}_{t}-\mathbf{U}\mathbf{J}\mathbf{I}_{P:S}\tilde{\mathbf{J}}^{-2}\bm{\eta}_{*},$ (43)
since $\varphi$ is an affine function with factor $\mathbf{J}$. Finally, Eqn. 43 boils down to:
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\mathbf{U}\mathbf{I}_{P:S}\bm{\eta}_{*},$ (44)
which is exactly Method 1 in parameterization $\bm{\theta}$.
Appendix F Convergence rate in a simple case
We study the convergence of the method presented in Section 4.1 (without anisotropic Nesterov’s cubic regularization):
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t},$ $\bm{\eta}_{t}$ $:=\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t},$ (45)

### v4

@@ Appendix 0.C Link with Cauchy’s steepest descent and Newton’s method
$\boldsymbol{\theta}_{t+1}:=\boldsymbol{\theta}_{t}-\mathbf{G}_{t}\mathds{1}_{S}(\mathds{1}_{S}^{T}\mathbf{G}_{t}\mathbf{H}_{t}\mathbf{G}_{t}\mathds{1}_{S})^{-1}\mathds{1}_{S}^{T}\mathbf{G}_{t}\mathbf{g}_{t}=\boldsymbol{\theta}_{t}-\mathbf{g}_{t}\frac{\mathbf{g}_{t}^{T}\mathbf{g}_{t}}{\mathbf{g}_{t}^{T}\mathbf{H}_{t}\mathbf{g}_{t}},$ (31)
@@
$\boldsymbol{\theta}_{t+1}$ $:=\boldsymbol{\theta}_{t}-\mathbf{G}_{t}(\mathbf{G}_{t}\mathbf{H}_{t}\mathbf{G}_{t})^{-1}\mathbf{G}_{t}\mathbf{g}_{t}.$ (32)
@@
$\boldsymbol{\theta}_{t+1}$ $:=\boldsymbol{\theta}_{t}-\mathbf{G}_{t}\mathbf{x}=\boldsymbol{\theta}_{t}-\mathbf{H}_{t}^{-1}\mathbf{g}_{t},$ (34)
@@
Appendix 0.D Anisotropic Nesterov cubic regularization
@@
$T(\boldsymbol{\eta}):=-\boldsymbol{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\boldsymbol{\eta}\bar{\mathbf{H}}\boldsymbol{\eta}+\frac{\lambda_{\mathrm{int}}}{6}\|\mathbf{D}\boldsymbol{\eta}\|^{3}.$ (35)
The function $T$ is strictly convex if, and only if, $\bar{\mathbf{H}}$ is positive definite. Moreover, $T$ is differentiable twice and has at least one global minimum $\boldsymbol{\eta}_{*}$, so $\frac{\mathrm{d}T}{\mathrm{d}\boldsymbol{\eta}}(\boldsymbol{\eta}_{*})=0$. Therefore, we first look for the solutions of the equation $\frac{\mathrm{d}T}{\mathrm{d}\boldsymbol{\eta}}(\boldsymbol{\eta})=0$.
@@
$\frac{\mathrm{d}T}{\mathrm{d}\boldsymbol{\eta}}(\boldsymbol{\eta})$ $=-\bar{\mathbf{g}}+\bar{\mathbf{H}}\boldsymbol{\eta}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\boldsymbol{\eta}\|\mathbf{D}^{2}\boldsymbol{\eta}$
$=-\bar{\mathbf{g}}+\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\boldsymbol{\eta}\|\mathbf{D}^{2}\right)\boldsymbol{\eta},$
@@
$\bar{\mathbf{g}}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\boldsymbol{\eta}\|\mathbf{D}^{2}\right)\boldsymbol{\eta}.$ (36)
Let $\boldsymbol{\eta}^{\prime}:=\mathbf{D}\boldsymbol{\eta}$. Eqn. (36) is then equivalent to:
$\bar{\mathbf{g}}$ $=\left(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}\|\boldsymbol{\eta}^{\prime}\|\mathbf{D}\right)\boldsymbol{\eta}^{\prime}.$
$=\frac{\lambda_{\mathrm{int}}}{2}\mathbf{D}\left(\frac{2}{\lambda_{\mathrm{int}}}\mathbf{D}^{-1}\bar{\mathbf{H}}\mathbf{D}^{-1}+\|\boldsymbol{\eta}^{\prime}\|\mathbf{I}\right)\boldsymbol{\eta}^{\prime}$
@@
$\bar{\mathbf{g}}$ $=\frac{\lambda_{\mathrm{int}}}{2}\mathbf{D}\left(\mathbf{K}+\|\boldsymbol{\eta}^{\prime}\|\mathbf{I}\right)\boldsymbol{\eta}^{\prime}$ (37)
@@
In this case, Eqn. (37) is equivalent to:
$\boldsymbol{\eta}^{\prime}$ $=\frac{2}{\lambda_{\mathrm{int}}}\left(\mathbf{K}+\|\boldsymbol{\eta}^{\prime}\|\mathbf{I}\right)^{-1}\mathbf{D}^{-1}\bar{\mathbf{g}}.$
Now, let $r=\|\boldsymbol{\eta}^{\prime}\|$. We want to solve:
@@
Trivially: $\boldsymbol{\eta}$ solution of (36) $\Rightarrow$ $\mathbf{D}\boldsymbol{\eta}$ solution of (37) $\Rightarrow$ $\|\mathbf{D}\boldsymbol{\eta}\|$ solution of (38). Reciprocally: $r$ solution of (38) $\Rightarrow$ $\boldsymbol{\eta}^{\prime}:=(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}r\mathbf{D})^{-1}\bar{\mathbf{g}}$ solution of (37) $\Rightarrow$ $\mathbf{D}^{-1}\boldsymbol{\eta}^{\prime}$ solution of (36).
Therefore, in order to find the unique global minimum of $T$, it is sufficient to solve Eqn. (38). This is doable numerically.
@@
We follow the procedure proposed in [24, Section 5]. Let $\lambda_{\mathrm{min}}$ be the minimum eigenvalue of $\mathbf{K}$. So, $\lambda_{\mathrm{min}}\leq 0$. Following [24], we look for the unique $\boldsymbol{\eta}^{\prime}$ belonging to $\mathcal{C}:=\{\boldsymbol{\eta}^{\prime}\in\mathbb{R}^{S}:\|\boldsymbol{\eta}^{\prime}\|>|\lambda_{\mathrm{min}}|\}$, which is also the solution of maximum norm of Eqn. (37). Conditionally to $\boldsymbol{\eta}^{\prime}\in\mathcal{C}$, $(\mathbf{K}+\|\boldsymbol{\eta}^{\prime}\|\mathbf{I})$ is invertible. So we only need to solve:
@@
which has exactly one solution $r_{*}$. Then, we compute $\boldsymbol{\eta}_{*}:=\mathbf{D}^{-1}(\bar{\mathbf{H}}\mathbf{D}^{-1}+\frac{\lambda_{\mathrm{int}}}{2}r_{*}\mathbf{D})^{-1}\bar{\mathbf{g}}$.
Appendix 0.E Invariance by subset-wise affine reparameterization
0.E.1 Motivation
The choice of the best per-layer parameterization is still a debated question. On the theoretical side, the standard parameterization cannot be used to train very wide networks, because it leads to a diverging first gradient step [33]. Besides, the NTK parameterization is widely used in theoretical works in order to manage the infinite-width limit [14, 8, 2, 17, 21]. On the practical side, the standard parameterization is preferred over the NTK one because it leads to better results, both in terms of training and generalization.
@@ 0.E.1 Motivation
0.E.2 Claim
We consider a parameter $\tilde{\boldsymbol{\theta}}$ such that $\boldsymbol{\theta}=\varphi(\tilde{\boldsymbol{\theta}})$, where $\varphi$ is an invertible map, affine on each subset of parameters. Therefore, its Jacobian is: $\mathbf{J}=\mathrm{Diag}(\alpha_{1},\cdots,\alpha_{p})$, where, for all $1\leq s\leq S$ and $1\leq p_{1},p_{2}\leq P$, we have:
@@ 0.E.2 Claim
We want to compare the training trajectory of $\mathcal{L}(\boldsymbol{\theta})$ and $\mathcal{L}(\varphi(\tilde{\boldsymbol{\theta}}))$ when using Method 1. For any quantity $\mathbf{x}$ computed with the parameterization $\boldsymbol{\theta}$, we denote by $\tilde{\mathbf{x}}$ its counterpart computed with the parameterization $\tilde{\boldsymbol{\theta}}$.
We compute $\tilde{\boldsymbol{\eta}}_{*}$. Equation (10) gives:
$\tilde{\boldsymbol{\eta}}_{*}=\left(\tilde{\bar{\mathbf{H}}}+\frac{\lambda_{\mathrm{int}}}{2}\|\tilde{\mathbf{D}}\tilde{\boldsymbol{\eta}}_{*}\|\tilde{\mathbf{D}}^{2}\right)^{-1}\tilde{\bar{\mathbf{g}}}.$ (41)
@@
since $\mathbf{J}$ and $\mathbf{U}$ are diagonal. And, since $\mathbf{D}_{ii}=|(\mathbf{D}_{\boldsymbol{\theta}}^{(3)}(\mathbf{u}))_{iii}|^{1/3}$, then $\tilde{\mathbf{D}}_{ii}=a_{i}^{2}\mathbf{D}_{ii}$, thus $\tilde{\mathbf{D}}=\tilde{\mathbf{J}}^{2}\mathbf{D}$.
Thus, Eqn. (41) becomes:
$\tilde{\boldsymbol{\eta}}_{*}=\left(\tilde{\mathbf{J}}^{2}\bar{\mathbf{H}}\tilde{\mathbf{J}}^{2}+\frac{\lambda_{\mathrm{int}}}{2}\|\tilde{\mathbf{J}}^{2}\mathbf{D}\tilde{\boldsymbol{\eta}}_{*}\|\tilde{\mathbf{J}}^{4}\mathbf{D}^{2}\right)^{-1}\tilde{\mathbf{J}}^{2}\bar{\mathbf{g}},$
@@
$\tilde{\mathbf{J}}^{2}\tilde{\boldsymbol{\eta}}_{*}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\tilde{\mathbf{J}}^{2}\tilde{\boldsymbol{\eta}}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}}.$
Therefore, $\tilde{\boldsymbol{\eta}}_{*}$ is a solution of Eqn. (10) in the parameterization $\tilde{\boldsymbol{\theta}}$ if, and only if, $\tilde{\mathbf{J}}^{2}\tilde{\boldsymbol{\eta}}_{*}$ is a solution in the parameterization $\boldsymbol{\theta}$. Moreover, $\|\tilde{\mathbf{D}}\tilde{\boldsymbol{\eta}}_{*}\|=\|\mathbf{D}\mathbf{J}^{2}\tilde{\boldsymbol{\eta}}_{*}\|$, so $\tilde{\boldsymbol{\eta}}_{*}$ is the solution of maximum norm $\|\tilde{\mathbf{D}}\tilde{\boldsymbol{\eta}}_{*}\|$ of (10) with parameterization $\tilde{\boldsymbol{\theta}}$ iff $\tilde{\mathbf{J}}^{2}\tilde{\boldsymbol{\eta}}_{*}$ is a the solution of maximum norm $\|\mathbf{D}\mathbf{J}^{2}\tilde{\boldsymbol{\eta}}_{*}\|$ of (10) with parameterization $\boldsymbol{\theta}$.
Thus, $\boldsymbol{\eta}_{*}=\tilde{\mathbf{J}}^{2}\tilde{\boldsymbol{\eta}}_{*}$, and the update step in parameterization $\tilde{\boldsymbol{\theta}}$ is:
$\tilde{\boldsymbol{\theta}}_{t+1}$ $=\tilde{\boldsymbol{\theta}}_{t}-\tilde{\mathbf{U}}_{t}\mathbf{I}_{P:S}\tilde{\boldsymbol{\eta}}_{*}$
$=\tilde{\boldsymbol{\theta}}_{t}-\tilde{\mathbf{U}}_{t}\mathbf{I}_{P:S}\tilde{\mathbf{J}}^{-2}\boldsymbol{\eta}_{*},$
@@
$\mathbf{J}^{-1}\boldsymbol{\theta}_{t+1}$ $=\mathbf{J}^{-1}\boldsymbol{\theta}_{t}-\mathbf{U}\mathbf{J}\mathbf{I}_{P:S}\tilde{\mathbf{J}}^{-2}\boldsymbol{\eta}_{*},$ (42)
since $\varphi$ is an affine function with factor $\mathbf{J}$. Finally, Eqn. 42 boils down to:
$\boldsymbol{\theta}_{t+1}$ $=\boldsymbol{\theta}_{t}-\mathbf{U}\mathbf{I}_{P:S}\boldsymbol{\eta}_{*},$ (43)
which is exactly Method 1 in parameterization $\boldsymbol{\theta}$.
Appendix 0.F Convergence rate in a simple case
We study the convergence of the method presented in Section 4 (without anisotropic Nesterov’s cubic regularization):
$\boldsymbol{\theta}_{t+1}$ $=\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t},$ $\boldsymbol{\eta}_{t}$ $:=\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t},$ (44)